# Enhanced sensitivity to trace 238U impurity of sapphire via coincidence neutron activation analysis

### v1

@@ IV.4 Analysis results
Average $0.947\pm 0.061$ $002\pm 0.074 1.1$$31\pm 0.070 1.1$$180\pm 0.098$
Table 1: Elemental concentrations determined for sapphire. Values for Mo, marked with a , are multiplied by a factor $10^{-6}$. Data reported in the single-$\gamma$ column are the weighted average of the Ge-A and Ge-B results. Concentrations marked with “FA” were obtained using the “fly ash” method; all others rely on the detector response model derived efficiency. Concentrations in the $\gamma$-$\gamma$ coincidence column were obtained using the two independent permutations of the coincidence condition and combining their weighted average. In addition, the results of the two independent $\rm{}^{239}Np$ cascades were combined by means of a weighted average. The Th and U concentration limits, in parenthesis, are based on the Feldman-Cousins method [33]. Errors are reported as the quadratic sum of the statistical and the 10% systematic error associated with the detector response model. For $\rm{}^{233}Pa$ and $\rm{}^{239}Np$, the errors are dominantly statistical. Concentrations reported as “UA 2017” were obtained using a different neutron flux and energy distribution than “UA 2024.” Various ratios are stated with their uncertainties to demonstrate analysis consistency between TUNL and UA.
@@
Figure 6: Time dependence of the $^{239}\text{Np}$ 277.6 keV $\upgamma\text{-ray}$ full absorption peak count rate as observed by Ge-A, requiring coincidence with Ge-B. Top: fly ash, bottom: sapphire. Red line: exponential decay model with the half-life fixed to that of $\rm{}^{239}Np$.
@@
The application of this method allowed us to reach ppt sensitivity for $\rm{}^{238}U$. We consider this result to be an experimental demonstration of the computational projection published in Ref. [17].
For $\rm{}^{232}Th$ contained in the sapphire, only single $\upgamma\text{ counting}$ could be used. However, compared to the 2017 analysis, two detectors were available rather than one. Also, here no evidence of contamination is found. We report the analysis result and Feldman-Cousins limit in parenthesis.
@@
Table 2: Compilation of $\rm{}^{232}Th$ and $\rm{}^{238}U$ radioassay results found in the literature and derived from our work. In some cases constraints on $\rm{}^{26}Al$, $\rm{}^{40}K$, $\rm{}^{60}Co$ and $\rm{}^{137}Cs$ activities exist but are not shown. The assay method determines which part of the decay series has been probed. In the case of NAA, $\rm{}^{232}Th$ and $\rm{}^{238}U$ are measured. In case of Ge counting (Ge), lower chain members are counted and their activities have been converted into concentrations of the chain heads by assuming secular equilibrium. For nEXO data, limits were determined using the Feldman-Cousins approach [33]. The $\rm{}^{238}U$ constraint marked with a is an exception. The raw result is 2.4 standard deviations negative, resulting in an inflated sensitivity of $\rm<1.8\;ppt$. We decided to not interpret this result as a statistical fluctuation and state the result using the flip-flop method instead, not taking credit for the unphysicality. Crackle and alumina are precursors to the sapphire crystal growth.
To provide a context for our results, we are comparing them to other radioactivity measurements found in the literature. Table 2 lists all radioactivity measurements we could find. The results attributed to nEXO will appear in a paper that is in preparation. As one can see from the Table 2, there is broad interest in the radioactivity content of this material.
@@ Acknowledgements.
The authors thank the nEXO collaboration for many stimulating and useful discussions. This work was supported in part by the U.S. Department of Energy, Office of Nuclear Physics, under Grant No. DE-FG02-01ER41166,  DE-FG02-97ER41033, DE-SC0023053, DE-SC0024666, the Laboratory Directed Research and Development and Early Career Research program at Pacific Northwest National Laboratory (PNNL), a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830, and by the U.S. National Science Foundation under award No. NSF-PHY2111213.

### v3

@@ IV.4 Analysis results
Table 1: Elemental concentrations determined for Saint-Gobain G3 sapphire. Values for Mo, marked with a , are multiplied by a factor $10^{-6}$. Data reported in the single-$\gamma$ column are the weighted average of the Ge-A and Ge-B results. The reported uncertainties correspond to that of the weighted average, discussed in section IV.2. Concentrations marked with “FA” were obtained using the “fly ash” method; all others rely on the detector response-model derived efficiency. Concentrations in the $\gamma$-$\gamma$ coincidence column were obtained using the two independent permutations of the coincidence condition and combining their weighted average (see section IV.2). In addition, the results of the two independent $\rm{}^{239}Np$ cascades were combined by means of a weighted average. The Th and U concentration limits, in parenthesis, are based on the Feldman-Cousins method [21]. Errors are reported as the quadratic sum of the statistical and the 10% systematic error associated with the detector response model. For $\rm{}^{233}Pa$ and $\rm{}^{239}Np$, the errors are dominantly statistical. Concentrations reported as “UA 2017” were obtained using a different neutron flux and energy distribution than “UA 2024.” Various ratios are stated with their uncertainties to demonstrate analysis consistency between TUNL and UA, different reactor neutron energy distributions, and singles and coincidence counting. The line denoted “Average” summarizes this comparison. The horizontal divider delineates the difference between analysis consistency checks and reporting of method-dependent sensitivity progress for $^{232}\text{Th}$ and $^{238}\text{U}$.
@@
Figure 7: Time dependence of the $^{239}\text{Np}$ 277.6 keV $\upgamma\text{-ray}$ full absorption peak count rate as observed by Ge-A, requiring coincidence with Ge-B. Top: fly ash, bottom: sapphire. Red line: exponential decay model with the half-life fixed to that of $\rm{}^{239}Np$. Rates with statistical errors obtained from the fit are shown.
@@
The application of the $\upgamma\upgamma\text{ coincidence counting}$ method allowed us to reach ppt sensitivity for $\rm{}^{238}U$. We consider this result to be an experimental demonstration of the computational projection published in Ref. [34].
For $\rm{}^{232}Th$ contained in the sapphire, only single $\upgamma\text{ counting}$ could be used. However, compared to the 2017 analysis, two detectors were available rather than one. Also, here no evidence of contamination is found. We report the analysis result and Feldman-Cousins limit in parenthesis. Fits to the single $\gamma$-spectra of the unobserved 106, 228, and 278 keV peaks in the 2024 TUNL sapphire data resulted in negative rates for all three peaks and both detectors. This behavior is not indicative of a statistical fluctuation. Consequently, we do not derive concentration values from that data.
@@
Table 2: Compilation of $\rm{}^{232}Th$ and $\rm{}^{238}U$ radioassay results found in the literature for sapphire and derived from our work. In some cases constraints on $\rm{}^{26}Al$, $\rm{}^{40}K$, $\rm{}^{60}Co$ and $\rm{}^{137}Cs$ activities exist but are not shown. Crackle and alumina are precursors to the sapphire crystal growth. The assay method determines which part of the decay series has been probed. In the case of NAA, $\rm{}^{232}Th$ and $\rm{}^{238}U$ are measured. In case of Ge counting (Ge), lower chain members are counted and their activities have been converted into concentrations of the chain heads by assuming secular equilibrium. For nEXO data, limits were determined using the Feldman-Cousins approach [21]. The $\rm{}^{238}U$ constraint marked with a is an exception. The raw result is 2.4 standard deviations negative, resulting in an inflated sensitivity of $\rm<1.8\;ppt$. We decided to not interpret this result as a statistical fluctuation and state the result using the flip-flop method instead, not taking credit for the unphysicality.
To provide a context for our results, we are comparing them to other sapphire radioactivity measurements found in the literature. Table 2 lists all radioactivity measurements we could find. The results attributed to nEXO will appear in a paper that is in preparation. As can be seen in Table 2, there is a broad interest in the radioactivity content of this material.
@@ Acknowledgements.
The authors thank the nEXO collaboration for many stimulating and useful discussions. This work was supported in part by the U.S. Department of Energy Office of Nuclear Physics and Office of High Energy Physics under Grant No. DE-FG02-01ER41166, DE-SC0012447,  DE-FG02-97ER41033, DE-SC0023053, DE-SC0024666, the Laboratory Directed Research and Development and Early Career Research program at Pacific Northwest National Laboratory (PNNL), a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830, and by the U.S. National Science Foundation under award No. NSF-PHY2111213.